OpenStudy (anonymous):

6^-3 What is the fraction of this?

Nnesha (nnesha):

bec 6 have negative exponent so you have to move to the bottom to take positive exponent for example \[\huge\ x^{-2} = \frac{ 1 }{ x^2 }\]

OpenStudy (anonymous):

so \[So \frac{ 1 }{ 6^{-3} }... which gives me 216... \]

Nnesha (nnesha):

nope not -3 when you move that 6 to the bottom exponent going to be positive

OpenStudy (alexandervonhumboldt2):

x^-a=1/x^a

Nnesha (nnesha):

did you understand??

OpenStudy (alexandervonhumboldt2):

1/(6*6*6)

OpenStudy (alexandervonhumboldt2):

=?

OpenStudy (anonymous):

okay so 1/216 would be the correct answer?

OpenStudy (alexandervonhumboldt2):

yeah

Nnesha (nnesha):

\[\ 6^{-3} = ????\]

OpenStudy (alexandervonhumboldt2):

i cant message you so i have to post it here you blocked me

Nnesha (nnesha):

well u got the answer so when move 6^-3 in at the bottom sign going to be  change and that's the reason we flip that bec we need positive exponent hope you understand good luck!